OpenStudy (anonymous):

Steve has the following scores on the exams for the semester: 89, 92, and 87. What will Steve need to make on the fourth exam to have a final average of an “A” (a 92 or better)? Use the Equation Editor to write your answer as an inequality.Answer

OpenStudy (anonymous):

\[(89 + 92 + 87 + m)\div 4 \ge 92\] so \[m \ge 100\]

OpenStudy (anonymous):

(89+92+87+x)/4=92 89+92+87+x =368 368-268=x=100